# The extinction of the contact process in a one-dimensional random environment with long-range interactions

### v1

@@ 2. Control of the environment, cut-points
Let $\{\tilde{\theta}_{j},\leavevmode\nobreak\ j\geqslant 1\}$ be a sequence of i.i.d. copies of $\tilde{\theta}$ and $G$ a geometric random variable with parameter $q$ independent of this sequence. In this way, we have the following equality in distribution
[…]
@@
$N_{M}=\big{\lvert}\{\langle x,y\rangle\colon\min\{x,y\}\in[[1,M]],\langle x,y\rangle\text{ is open}\}\big{\rvert}$
@@
It suffices to consider even values of $k$. Without loss of generality, we also assume that $k$ is positive. Let $\delta\in(2,s-2)$, and define $\varepsilon=(1+\delta)/(1+\alpha)-1$, where $\alpha=5\delta/6$.
@@
$N_{M}^{z}=\big{|}\big{\{}\langle x,y\rangle\in\mathcal{E}\colon\min\{x,y\}\in[[z+1,z+M]]\big{\}}\big{|},$
be the number of open edges linking a vertex in $[[z+1,z+M]]$ to another vertex lying to its right. Conditioning on $\{v_{k/2}=z\}$ and using Lemma 7 we get, for $M$ sufficiently large (depending on $\delta$).
@@ 3. The renormalization step
Let $\Phi=(\phi_{k})_{k\in{\mathbb{Z}}}$ be a sequence of independent positive random variables that we call a random environment. Conditional on $\Phi$, we denote by $P_{\Phi,p}$ the law of a site percolation process on $\mathbb{Z}\times\mathbb{Z}_{+}$ under which every site $(i,j)$ is open independently with probability $p^{\phi_{i}}$. In what follows, we will be interested in the setting where $\phi_{k}=|\mathcal{E}_{k}|$ for every $k$, because by (20) $P_{\Phi,p}$ is dominated stochastically by the law of the random process $(\mathbf{1}_{\text{$(k,j)$ is good}})_{(k,j)\in\mathbb{Z}\times\mathbb{Z}_{+}}$.
Since $(v_{k})_{k\in\mathbb{Z}\setminus\{0\}}$ are cut-points there are no open edges connecting $V_{k}$ to $V_{\ell}$ if $k\neq\ell$. Therefore, the existence of a semi-circuit of good vertices around the origin in $\mathbb{Z}\times\mathbb{Z}_{+}$, ensures that the contact process dies out.
@@
Figure 3. The renormalization scheme. We represent the event $\{(2k,2)\text{ is good}\}$. In the corresponding rectangle $[[v_{k}+1,v_{k+1}-1]]\times[T,2T]$ every site has a recovery mark (lozenge) and no edge has an infection attempt (double arrow).
@@ 4. Crossing events
A subset $R=[[a,b]]\times[[c,d]]\subset{\mathbb{Z}}\times{\mathbb{Z}}_{+}$ where $a<b$ and $0\leqslant c<d$ are integer numbers is called a rectangle. We will be interested in the occurrence of horizontal and vertical crossings spanning $R$:
[…]
The next result compares the crossing probabilities of a rectangle $R$ and a rescaled version of it under the site percolation measure $P_{\Phi,p}$ introduced in the previous section and the bond percolation measure $\mathbb{P}^{\Lambda}_{\rho}$, respectively. The content of the following lemma is based on a standard coupling of site and bond percolation and is reminiscent of [Grimm, Theorem 1.33].
@@
(21) $P_{\Phi,p}\big{(}\mathcal{C}_{h}([[2a+1,2b+1]]\times[[c,d]])\big{)}$$\geqslant$ $\mathbb{P}_{\rho}^{\Lambda}\big{(}\mathcal{C}_{h}([[a,b]]\times[[c,d]])\big{)},$
(22) $P_{\Phi,p}\big{(}\mathcal{C}_{v}([[2a+1,2b+1]]\times[[c,d]])\big{)},$$\geqslant$ $\mathbb{P}_{\rho}^{\Lambda}\big{(}\mathcal{C}_{v}([[a,b]]\times[[c,d]])\big{)}.$
@@
Let $R=[[2a+1,2b+1]]\times[[c,d]]$. Write $e_{1}=\langle(0,0),(2,0)\rangle$ and $e_{2}=\langle(0,0),(0,1)\rangle$ and define the following set of unoriented edges
$e(R)=\big{\{}\big{\langle}(2i+1,j),(2i+1,j)+e_{\ell}\big{\rangle}\colon i\in[[a,b]],j\in[[c,d]],\ell=1,2\big{\}}.$
@@
$(A_{0},B_{0})=\big{(}\{2a+1\}\times[[c,d]],\varnothing\big{)}.$
@@
$P_{\Phi,p}\big{(}\mathcal{C}_{h}(R)\big{)}$ $\geqslant\mathbb{P}\left(A_{\infty}\cap\big{(}\{2b+1\}\times[[c,d]]\big{)}\neq\varnothing\right)$
$=\mathbb{P}_{\rho}^{\Lambda}\big{(}\mathcal{C}_{h}([[a,b]]\times[[c,d]])\big{)},$
where, we recall, $\Lambda=\{\ldots,\phi_{-2}+1,\phi_{0}+1,\phi_{2}+1,\phi_{4}+1,\ldots\}$ and $\mathbb{P}^{\Lambda}_{\rho}$ is the quenched law governing the percolation process defined in (1). In the right-hand side we abuse notation and write $\mathcal{C}_{h}\big{(}([[a,b]]\times[[c,d]])\big{)}$ for the event that the respective rectangle is crossed horizontally by a path of open bonds. The rectangle $[[a,b]]\times[[c,d]]$ appears instead of $R$, because we are using edges of length $2$ in our exploration process.
@@
$P_{\Phi,p}\big{(}\mathcal{C}_{v}(R)\big{)}\geqslant\mathbb{P}_{\rho}^{\Lambda}\big{(}\mathcal{C}_{v}([[a,b]]\times[[c,d]])\big{)}.$
@@ 5. Percolation on stretched lattices
(27) $\begin{cases}x_{k}=x_{k-1}+\xi_{k},\leavevmode\nobreak\ k\in\mathbb{Z}_{+},\\ x_{k}=x_{k+1}-\xi_{k+1},\leavevmode\nobreak\ k\in\mathbb{Z}_{-},\end{cases}$
@@
$V(\mathcal{L}_{\Lambda})$ ::=:= $\Lambda\times\mathbb{Z}_{+}$
$E(\mathcal{L}_{\Lambda})$ ::=:= $\big{\{}\langle(x_{i},n),(x_{j},m)\rangle\colon|i-j|+|n-m|=1\big{\}}.$

### v2

@@ 2. Control of the environment, cut-points
Let $\{\tilde{\theta}_{j},\penalty 10000\ j\geqslant 1\}$ be a sequence of i.i.d. copies of $\tilde{\theta}$ and $G$ a geometric random variable with parameter $q$ independent of this sequence. In this way, we have the following equality in distribution
$\begin{split}\eta_{1}&=\sum_{j=2i_{1}}^{2i_{2}-1}\theta_{j}=\theta_{i_{1}}+\theta_{2i_{2}-1}+\left[\sum_{j=i_{1}+1}^{i_{2}-1}(\theta_{2j-1}+\theta_{2j})\right]1_{\{i_{2}\neq i_{1}+1\}}\penalty 10000\ \\ &\overset{D}{=}\penalty 10000\ 2+\left[\sum_{j=1}^{G-1}\tilde{\theta}_{j}\right]1_{\{G\neq 1\}}.\end{split}$
@@
$N_{M}=\big\lvert\{\langle x,y\rangle\colon\min\{x,y\}\in[\hskip-1.00006pt[1,M]\hskip-1.00006pt],\langle x,y\rangle\text{ is open}\}\big\rvert$,
where the notation $[[a,b]]$ denotes the set $a,b]\cap{\mathbb{Z}}$.
@@
It suffices to consider even values of $k$. Without loss of generality, we also assume that $k$ is positive. Let $\delta\in(0,s-2)$, and define $\varepsilon=(1+\delta)/(1+\alpha)-1$, where $\alpha=5\delta/6$.
@@
$N_{M}^{z}=\big|\big\{\langle x,y\rangle\in\mathcal{E}\colon\min\{x,y\}\in[\hskip-1.00006pt[z+1,z+M]\hskip-1.00006pt]\big\}\big|,$
be the number of open edges linking a vertex in $\hskip-1.00006pt[z+1,z+M]\hskip-1.00006pt]$ to another vertex lying to its right. Conditioning on $\{v_{k/2}=z\}$ and using Lemma 7 we get, for $M$ sufficiently large (depending on $\delta$).
@@ 3. The renormalization step
Since $(v_{k})_{k\in\mathbb{Z}\setminus\{0\}}$ are cut-points there are no open edges connecting $V_{k}$ to $V_{\ell}$ if $|k-\ell|>1$. Consequently, if the infection enters a good box, it can only propagate to an adjacent box sharing a common face. However, this is impossible if the latter is also good. This implies that the existence of a semi-circuit of good vertices around the origin in $\mathbb{Z}\times\mathbb{Z}_{+}$, ensuring that the contact process dies out.
$v_{k-1}$$v_{k}$$v_{k+1}$$v_{k+2}$0TT2T2T
Figure 4. The renormalization scheme. We represent the event $\{(2k,1)\text{ is good}\}$. In the corresponding rectangle $\hskip-1.00006pt[v_{k}+1,v_{k+1}-1]\hskip-1.00006pt]\times[T,2T]$ every site has a recovery mark (lozenge) and no edge has an infection attempt (double arrow).
@@ 4. Crossing events
A subset $R=[\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]\subset{\mathbb{Z}}\times{\mathbb{Z}}_{+}$ where $a<b$ and $0\leqslant c<d$ are integer numbers is called a rectangle. We will be interested in the occurrence of horizontal and vertical crossings spanning $R$:
$\mathcal{C}_{h}(R)$ $=\big\{\{a\}\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]\overset{R}{\longleftrightarrow}\{b\}\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]\big\},$
$\mathcal{C}_{v}(R)$ $=\big\{[\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times\{c\}\overset{R}{\longleftrightarrow}[\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times\{d\}\big\}$
where $A\overset{S}{\longleftrightarrow}B$ stands for the existence of a path of open vertices connecting $A$ to $B$ while remaining inside $S$. We abuse notation and still write $\mathcal{C}_{h}(R)$ and $\mathcal{C}_{v}(R)$ for crossings of open edges in the rectangle $R$.
The next result compares the crossing probabilities of a rectangle $R$ and a rescaled version of it under the site percolation measure $P_{\Phi,p}$ introduced in the previous section and the bond percolation measure $\mathbb{P}^{\Lambda}_{\rho}$, respectively. The content of the following lemma is based on a standard coupling of site and bond percolation and is reminiscent of [11, Theorem 1.33].
@@
(21) $P_{\Phi,p}\big(\mathcal{C}_{h}([\hskip-1.00006pt[2a+1,2b+1]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big)$ $\geqslant$ $\mathbb{P}_{\rho}^{\Lambda}\big(\mathcal{C}_{h}([\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big),$
(22) $P_{\Phi,p}\big(\mathcal{C}_{v}([\hskip-1.00006pt[2a+1,2b+1]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big),$ $\geqslant$ $\mathbb{P}_{\rho}^{\Lambda}\big(\mathcal{C}_{v}([\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big).$
@@
Let $R=[\hskip-1.00006pt[2a+1,2b+1]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]$. Write $e_{1}=\langle(0,0),(2,0)\rangle$ and $e_{2}=\langle(0,0),(0,1)\rangle$ and define the following set of unoriented edges
$e(R)=\big\{\big\langle(2i+1,j),(2i+1,j)+e_{\ell}\big\rangle\colon i\in[\hskip-1.00006pt[a,b]\hskip-1.00006pt],j\in[\hskip-1.00006pt[c,d]\hskip-1.00006pt],\ell=1,2\big\}.$
@@
$(A_{0},B_{0})=\big(\{2a+1\}\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt],\varnothing\big).$
@@
$P_{\Phi,p}\big(\mathcal{C}_{h}(R)\big)$ $\geqslant\mathbb{P}\left(A_{\infty}\cap\big(\{2b+1\}\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]\big)\neq\varnothing\right)$
$=\mathbb{P}_{\rho}^{\Lambda}\big(\mathcal{C}_{h}([\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big),$
where, we recall, $\Lambda=\{\ldots,\phi_{-2}+1,\phi_{0}+1,\phi_{2}+1,\phi_{4}+1,\ldots\}$ and $\mathbb{P}^{\Lambda}_{\rho}$ is the quenched law governing the percolation process defined in (1). In the right-hand side we abuse notation and write $\mathcal{C}_{h}\big(([\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big)$ for the event that the respective rectangle is crossed horizontally by a path of open bonds. The rectangle $\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt]$ appears instead of $R$, because we are using edges of length $2$ in our exploration process.
@@
$P_{\Phi,p}\big(\mathcal{C}_{v}(R)\big)\geqslant\mathbb{P}_{\rho}^{\Lambda}\big(\mathcal{C}_{v}([\hskip-1.00006pt[a,b]\hskip-1.00006pt]\times[\hskip-1.00006pt[c,d]\hskip-1.00006pt])\big).$
@@ 5. Percolation on stretched lattices
(27) $\begin{cases}x_{k}=x_{k-1}+\xi_{k},\penalty 10000\ k\in\mathbb{Z}_{+},\\ x_{k}=x_{k+1}-\xi_{k+1},\penalty 10000\ k\in\mathbb{Z}_{-},\end{cases}$
@@
$V(\mathcal{L}_{\Lambda})$ $:=$ $\Lambda\times\mathbb{Z}_{+}$
$E(\mathcal{L}_{\Lambda})$ $:=$ $\big\{\langle(x_{i},n),(x_{j},m)\rangle\colon|i-j|+|n-m|=1\big\}.$
@@
Remark 5.
By the independence of the states of each edge in the long-range percolation graph $G=(V,\mathcal{E})$ the $|\mathcal{E}_{2k}|$ may assume any large enough integer value, so the aperiodicity assumption will not pose any serious restriction for our arguments.
Let us now fix two sequences $L_{0},L_{1},L_{2},\ldots$ and $H_{0},H_{1},H_{2},\ldots$ that will represent the horizontal length and vertical height scales along which we will analyze the model. They are defined recursively by fixing integers $L_{0}>1$, $H_{0}>1$ and constants $\gamma>1$ and $\mu\in(0,1)$ and then setting for every $k\geqslant 1$,
(32) $L_{k}=L_{k-1}\lfloor L_{k-1}^{\gamma-1}\rfloor,\qquad\textrm{ and }\qquad H_{k}=2\lceil\exp(L_{k}^{\mu})\rceil H_{k-1}.$
Remark 6.
For technical reasons, the initial scale $L_{0}$, has to be chosen sufficiently large, depending on $\varepsilon$. The reader is referred to [14, Conditions (i), (ii) and (iii), pp. 3153] for the exact requirements on $L_{0}$. In addition, the exponent $\gamma$ cannot be too large while $\mu$ has to be taken sufficiently close to one. The following choices
(33) $\gamma=1+\frac{\varepsilon}{2(\varepsilon+4)}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\textrm{ and }\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\mu=\frac{1}{2}\Big(1+\frac{1}{\gamma}\Big).$
will satisfy the requirements (see [14, Eq. (19) pp. 3153 and Eq.(32) pp. 3156]).
For $k\in\mathbb{Z}_{+}$, let us partition $\mathbb{R}_{+}$ into intervals of length $L_{k}$
$I_{j}^{k}:=\big[jL_{k},(j+1)L_{k}\big),\textrm{ with }j\in\mathbb{Z}.$
We will be interested in studying the occurrence of horizontal and vertical crossings of rectangles with length of order $L_{k}$ and height of order $H_{k}$.
For $i,j,k\in\mathbb{Z}_{+}$ denote
(34) $C_{i,j}^{k}:=\mathcal{C}_{h}\Big(\big(I^{k}_{i}\cup I^{k}_{i+1}\big)\times\big[jH_{k},(j+1)H_{k}\big)\Big)$
(35) $D_{i,j}^{k}:=\mathcal{C}_{v}\Big(I^{k}_{i}\times\big[jH_{k},(j+2)H_{k}\big)\Big).$
These events are illustrated in Figure 6.
$I_{i}^{k}$$I_{i+1}^{k}$$jH_{k}$$(j+1)H_{k}$$I_{i}^{k}$$jH_{k}$$(j+1)H_{k}$$(j+2)H_{k}$
Figure 6. Illustration of the events $C_{i,j}^{k}$ (left) and $D_{i,j}^{k}$ (right).
For $k\geqslant 1$, each $I_{j}^{k}$ can be partitioned into disjoint intervals from scale k1k-1:
(36) $I_{j}^{k}=\bigcup_{i\in l_{k,j}}I_{i}^{k-1},$
where the union runs over the set of indices
(37) $l_{k,j}=\big\{j\lfloor L_{k-1}^{\gamma-1}\rfloor,\cdots,(j+1)\lfloor L_{k-1}^{\gamma-1}\rfloor-1\big\}.$
Now fix an environment $\Lambda\subseteq\mathbb{Z}$. Intervals will be recursively classified either good or bad as follows.
  • at the bottom scale $k=0$, declare an interval bad if it does not intersect $\Lambda$ (see Figure 7);
  • at higher scales $k\geqslant 1$, declare an interval bad if it contains two non-consecutive bad intervals from scale k1k-1 (see Figure 8).
[…]
Figure 7. Bottom scale. Sites are colored black or white according to whether they belong to $\Lambda$ or not, respectively. Good intervals contain at least one point of $\Lambda$.
[…]
Figure 8. The configurations of good and bad intervals at scales $k$, $k+1$, and $k+2$. Good intervals are represented by thick solid lines, and bad intervals are represented by thin solid lines.
Let $\rho$ be a random variable with distribution:
(38) $\mathbb{P}(\rho=k):=\dfrac{1}{\mathbb{E}(\xi)}\sum_{i=k+1}^{\infty}\mathbb{P}(\xi=i),\ \textrm{for every }k\in\mathbb{Z}_{+},$
independent of everything else. Roughly, $\rho$ is the delay that should be introduced to the renewal process $\Lambda\cap\mathbb{Z}_{+}$ in order to make it stationary. In other words, modifying (27) to start with $x_{0}=\rho$ instead of $x_{0}=0$, the one-sided renewal process $\Lambda_{+}=\{x_{0},x_{1},x_{2},\ldots\}$ is stationary. Let us denote $\upsilon_{\xi}^{\rho,+}$ the law of that process. This allows us to define
(39) $p_{k}:=\upsilon^{\rho,+}_{\xi}\big(\text{$I_{0}^{k}$ is bad}\big)=\upsilon^{\rho,+}_{\xi}\big(\text{$I_{j}^{k}$ is bad}\big),$
where the second equality follows from the stationarity of $\rho$.
Let us define, for every $i,j,k\in\mathbb{Z}_{+}$ and $p\in(0,1)$,
(40) $q_{k}(p;i,j):=\max\Bigg\{\underset{\begin{subarray}{c}\Lambda;\ I^{k}_{i}\textrm{ and }\\ I^{k}_{i+1}\textrm{are good}\end{subarray}}{\max}\mathbb{P}^{\Lambda}_{p}\big((C_{i,j}^{k})^{c}\big),\underset{\begin{subarray}{c}\Lambda;\ I^{k}_{i}\textrm{ is}\\ \textrm{good}\end{subarray}}{\max}\mathbb{P}^{\Lambda}_{p}\big((D_{i,j}^{k})^{c}\big)\Bigg\}.$
Translation invariance allows us to write, for every $k\in\mathbb{Z}_{+}$,
(41) $q_{k}(p)\vcentcolon=q_{k}(p;0,0)=q_{k}(p;i,j),\textrm{ for any }i,j\in\mathbb{Z}_{+}.$
The next two lemmas, taken from [14] establish upper bounds for $p_{k}$ and $q_{k}(p)$.
Lemma 9 (Lemma 3.1 in [14]).
For every $k\in\mathbb{Z}_{+}$ we have
(42) $p_{k}\leqslant L_{k}^{-\varepsilon/2}.$
Lemma 10 (Lemma 3.2 in [14]).
There exist $c_{3}=c_{3}(L_{0},\varepsilon)\in\mathbb{Z}_{+}$ and $p=p(L_{0},\varepsilon,c_{3})$ sufficiently close to 1 such that
(43) $q_{k}(p)\leqslant\exp(-L_{k}^{\beta}),\,\,\,\,\textrm{ for any }k\geqslant c_{3},$
where the exponent is given by
(44) $\beta=1-\frac{\varepsilon}{8(\varepsilon+4)}.$
Remark 7.
That the value of $\beta$ given in (44) works, follows from our choices of $\gamma$ and $\mu$ in (33), together with [14, Eq. (37), pp. 3156].
Let us write $\upsilon_{\xi}^{n,+}$ for the one-sided renewal with interarrival distribution $\xi$ starting at position $n\geqslant 0$.
As a direct consequence of Lemma 9 we have
Lemma 11.
Let $\xi$ satisfy (30) and (31). Then, for every integer $n\geqslant 0$, we have
(45) $\upsilon_{\xi}^{n,+}\big(\text{$I^{k}_{0}$ and $I^{k}_{1}$ are good eventually}\big)=1.$
Proof.
By (42), we have
$\sum_{k=0}^{\infty}\upsilon^{\rho,+}_{\xi}\big(\text{$I^{k}_{0}$ or $I^{k}_{1}$ are bad}\big)\leqslant\sum_{k=0}^{\infty}L_{k}^{-\varepsilon/2}<\infty.$
Therefore,
$\upsilon^{\rho,+}_{\xi}\big(\text{$I^{k}_{0}$ and $I^{k}_{1}$ are good eventually}\big)=1.$
By (38), $\rho$ assumes any positive integer value with positive probability. Hence,
[…]
Lemma 12.
Let $\Lambda=\big\{\xi_{k}\big\}_{k\in\mathbb{Z}}$ where $\xi_{k}=|\mathcal{E}_{2k}|+1$. Then, for almost every realization of $\Lambda$, we have
(46) $\text{$I^{k}_{-2}$, $I^{k}_{-1}$, $I^{k}_{0}$ and $I^{k}_{1}$ are good eventually}.$
Proof.
Let us write $\xi$ for the law of $\xi_{k}$ with $k\neq 0$. We also write $\bar{\upsilon}_{\xi}^{0}$ for the law of the two-sided renewal process similar to $\upsilon_{\xi}^{0}$, with the exception that the distribution of $\xi_{0}$ is that of $|\mathcal{E}_{0}|+1$. Recall that $\upsilon_{\xi}^{n,+}$ denotes the one-sided renewal with interarrival distribution $\xi$ starting at position $n\geqslant 0$. We have
[…]
where we have used independence of what occurs on both sides of the origin and reflection invariance of the one-sided process.
By (45) with $n=0$, $\upsilon_{\xi}^{0,+}\big(\text{$I^{k}_{0}$ and $I^{k}_{1}$ are good eventually}\big)=1$, so it suffices to prove that
${\upsilon}_{\xi}^{|\mathcal{E}_{0}|+1,+}\big(\text{$I^{k}_{0}$ and $I^{k}_{1}$ are good eventually}\big)=1.$
In fact we have,
[…]
where we have also used (45) at the last step. ∎
Let us now consider the events
(49) $F_{k}=\Big(\bigcap_{i={-2}}^{0}C^{k}_{i,1}\Big)\cap\Big(\bigcap_{i={-2}}^{1}D^{k}_{i,0}\Big).$
See Figure 9.
[…]
Figure 9. Occurrence of the crossing events $C_{i,1}^{k}$, i=2,1,0i=-2,-1,0 together with $D_{i,0}^{k}$ i=2,1,0,1i=-2,-1,0,1 entails the existence of a dual circuit about the origin.
Lemma 13.
Let $\Lambda=\big\{|\mathcal{E}_{2k}|+1\big\}_{k\in\mathbb{Z}}$. Then, for every $p$ sufficiently large,
(50) $\mathbb{P}^{\Lambda}_{p}(F_{k}\text{ eventually})=1,$
for almost all realizations of $\Lambda$.
Proof.
Conditional on a realization of $\Lambda$ for which $I^{k}_{-2}$, $I^{k}_{-1}$, $I^{k}_{0}$ and $I^{k}_{1}$ are good eventually Lemma 10 yields
(51) $\sum_{k\geqslant k_{0}}\mathbb{P}_{p}^{\Lambda}(F_{k}^{c})\leqslant\sum_{k\geqslant k_{0}}7\exp(-L_{k}^{\beta})<\infty.$
Hence, a Borel-Cantelli argument implies (50).
6. Proof of Theorem 1
We are now in a position to prove Theorem 1. As argued in the last paragraph of Section 3, it suffices to show that for a sufficiently small infection rate $\lambda$, there almost surely exists a semi-circuit of good vertices enclosing the origin in the renormalized lattice $\mathbb{Z}\times\mathbb{Z}_{+}$. Lemma 13 enables us to study the existence of such circuits in a different bond percolation process. We then apply Lemma 8 to compare the probabilities of crossing events, and thus of existence of semi-circuits, in both models. This lemma establishes that the probability of semi-circuits in the former percolation process is bounded below by the probability in the latter, with a certian parameter $\rho$ that can be made as large as we want by decreasing $\lambda$. As stated in Lemma 13, this lower bound equals 1, provided that $\rho$ is sufficiently large. The result follows.